Find McLaurin Series for Sin(x)


1
Expert's answer
2021-09-01T11:27:03-0400

"f(x)=sin(x),\;f(0)=0."

"f'(x)=cos(x),\;f'(0)=1."

"f''(x)=-sin(x),\;f''(0)=0."

"f'''(x)=-cos(x),\;f'''(0)=-1."

"f^{(4)}(x)=-sin(x),\;f^{(4)}(0)=0."

Since "f^{(4)}=f(x)" , the cycle of {0, 1, 0, -1} repeats itself, which means that every derivative of even degree gives 0 and that every derivative of odd degree alternates between 1 and -1. So, we have

"f(x)=\Sigma^\infin_{n=0}\frac{f^{(n)}(0)}{n!}x^n=\Sigma^\infin_{n=0}(-1)^n\frac{x^{2n+1}}{(2n+1)!}"